MONT 106Q – Mathematical Thinking Second Writing Assignment

The second of the larger paper assignments this semester will be devoted to the Histories of Herodotus. It is also designed to establish a baseline for your level of mastery of an important form of writing sometimes called the argumentative essay. Note: “argumentative” here does not mean that you are trying to pick a fight through your writing(!) It means that you are making a a particular response to a question – a thesis – and then presenting appropriate evidence and reasoning to support your claim(s) and convince your reader that what you are saying is correct. Your personal feelings and opinions are not so important this time. Instead, your essay should show that you have done some careful reading, and some serious thinking, about the subject. You may, of course, disagree with ideas presented in class if you want. The point is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 4-5 pages in length (double-spaced, single-sided). Please submit via Google Docs or as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Friday, October 7.

Paper topic

As we ask in this year’s CHQ theme, how then shall we live when borders can serve to define us, but we may also wish to transcend them? Herodotus’ world seems to be especially full of borders between different ethnic groups, between different political units, between different customs, between right and wrong behaviors or ways of doing things and so forth. Pick one of the following sections of the Histories:

a. Chapters 1.131-1.133 – religion and customs of the Persians b. Chapters 2.2-2.3 – the “experiment of Psammetichos” c. Chapters 2.35-2.37 – customs of the Egyptians

or another passage of your choice. If you want to look at a passage we discussed extensively in class, please discuss your choice with Prof. Little to make sure that it is suitable and that you will not just be repeating points from what we talked about. Closely analyze how Herodotus works the theme of boundaries or borders into his presentation. Exactly what border or borders are involved in your passage? Are they historical, political, geographical, cultural, moral, or something else? In what ways do those borders define people or set people apart? Does Herodotus consider crossing or transcending those borders and would he say that would be good or desirable? Also, where is Herodotus’s information coming from in your section? Does he describe his sources or indicate how he learned about what he is saying? Can you say anything about the historical validity of his account? Finally, if you think your passage offers some insight about the question “how then shall we live?” in our CHQ theme, feel free to include a discussion of that too.

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. Clearly state which section of the Histories you are going to analyze first and briefly summarize what happens in that section. Please don’t start with overly general statements like “Herodotus was a Greek writer who lived in the 5th century BCE,” or “The Histories of Herodotus is one of the first examples of historical writing to have survived.” Assume your reader knows who Herodotus was, when he lived, and what his main contributions were. In other words, this is a paper presenting an analysis of one section of the Histories, not a “book report.” After identifying your passage, state your thesis responding to the main question above. The thesis should be clearly stated and well focused. The reader (me) should not need to guess what your point is.

3. In the following paragraphs, using sound reasoning, present your evidence for making the claim in your thesis. Do not contradict what you said earlier. Do not ignore evidence that contradicts your ideas. If, on further reflection, you find that you do not agree with what you said before, revise the thesis(!)

4. In this sort of writing, a “move” that is often very effective is to temporarily entertain an opposing conclusion as a possibility, then explain the evidence that leads you to reject that conclusion. Try it!

5. Your evidence should consist of passages from the Landmark Herodotus, and interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small details and mention the little things that support your point).

6. Do not consult any other sources beyond the Landmark Herodotus. You may use the translated text of the Histories, and the discussions in the Introduction or any of the Appendices that might be useful. But please do not consult other translations, other books, web sites, etc. If you use direct quotations from the Landmark translation of Herodotus’s own text, at the end place the book, chapter, and section to identify the source. Here’s an example of what I’m looking for:

Herodotus discusses three different explanations for why the Nile floods each year. He dismisses the first of these, namely that the floods are caused by the Etesian winds forcing the water back into the channel of the river, saying “often the Etesian winds do not blow, and the Nile floods at the same time anyway.” (Book 2, Chapter 20, Section 2, page 125).

If you use a direct quotation from the Introduction or one of the Appendices, identify the source by saying it comes from the Introduction, or the appropriate Appendix and give the page where your quotation is found.

7. The paper as a whole should be well-organized and “flow” from one point to the next. Do not repeat pieces of evidence unless there is a very good reason for doing so. Do not get ahead of yourself. Say one thing at a time. If you decide you want to add something, do not just tack it on somewhere. Instead, figure out where it fits in the overall scheme of your argument and put it where it belongs.

8. Your sentences should be clear, readable, and not overly wordy. Aim for a reasonably formal but lively style. I will not say you cannot use the first person, but if you do, there should be a good reason for doing so (for example if you are explicitly saying your opinion is different from another possible interpretation). Careless or colloquial ways of saying things that are acceptable in casual conversation are not appropriate in this kind of writing. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.